the problem is 4n^2+25=-20x. What is the value of the discriminant and please state the number and typs of roots/solutions/zeros? Please show your work as well. I think that I following the following a + b+ c+= a is 4^2+b which is 20+c which is 25=0. I am so confused I am not even sure if that is right. If you could do this and show all work it is greatly appreciated.

Answer by stanbon(75887)

the problem is 4n^2+25=-20x. What is the value of the discriminant and please state the number and typs of roots/solutions/zeros?
-----
Rearrange:
4x^2 +20x + 25 = 0
-----
a = 4 ; b = 20 ; c = 25
------
Discriminant: b^2-4ac
-----
20^2 - 4*4*25 = 400 - 400 = 0
-----
The quadratic has two equal Real Number solutions.
==================================================
Cheers,
Stan H.
============